OpenStudy (anonymous):

Verify SEC(X) CSC(X) ----------- = COT(X) TAN(X) Please Help SIN(X) COS(X) Explain Steps

OpenStudy (anonymous):

\[\frac{ SEC(X) CSC(X) }{ SIN(X)COS(X) }=COT(X)TAN(X)\] Verify This Explain

OpenStudy (anonymous):

This is the solution by how did it get from line 1 to line 2

OpenStudy (anonymous):

\[\frac{ 1 }{ SIN(X)COS(X) } \to \frac{ COS(X)^{2} + SIN(X)^{2}}{ SIN(X)COS(X)}\]

OpenStudy (anonymous):

The question is wrong. The RHS gives 1 whereas the LHS is 1/(sinx*cosx)^2

OpenStudy (anonymous):

@mark_o. Can you help me again?

OpenStudy (anonymous):

i don;t understand how is the question wrong?

OpenStudy (anonymous):

Because the LHS is not equal to the RHS. I think the question is sec(x)cos(x)/sin(x)csc(x) =tan(x)cot(c)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i moved this http://openstudy.com/updates/509c8db9e4b0a825f21b1cf9